# Dynamic Rank Adjustment for Accurate and Efficient Neural Network Training

### v2

@@ 5.3 Computational Cost Analysis
Here, we analyze the computational overhead of dynamic-rank training, which is shown in Comp. columns in Table 3 and 6. First, the full-rank training cost is calculated as
@@
where $E$ is the total number of training epochs and $d$ is the model size. Then, the low-rank training cost is calculated as
@@
Thus, if $\rho$ is the same, $T_{DR}$ should be higher than $T_{SVD}$ and lower than $T_{F}$. For example, ResNet20 contains 272,762 trainable parameters and the reparameterized model with SVD ($\rho=0.5$) has 155,170 parameters. When $E=150$, $I=60$, and $D=135$, where $\phi=0.5$, $T_{DR}/T_{F}$ becomes $0.7844$, which is shown in Comp. column in Table 3. Our empirical study demonstrates that, by carefully tuning $I$ and $D$, a relatively small DID-I value provides comparable accuracy to the full-rank training. Small DID-I settings in dynamic-rank training makes its training cost more similar to that of the conventional low-rank training.
@@ 6 Conclusion
In this study, we propose a dynamic-rank training framework that restores the effective rank of model weights while preserving the efficiency of low-rank training. We also present two key insights for maximizing rank recovery by strategically interleaving full-rank epochs within low-rank training. Our extensive empirical study proves that scheduling full-rank epochs at the end of the high-noise regime and the beginning of the low-noise regime maximizes recovery of the model’s effective rank, thereby improving accuracy. The proposed dynamic-rank training scheme is general and readily applicable to any deep learning applications. We believe that harmonizing the proposed dynamic-rank training and other compute-efficient neural network training methods can be a promising future work.

### v3

@@ 5.2 Computational Cost Analysis
Here, we analyze the computational overhead of dynamic-rank training, which is shown in Comp. columns in Table 4 and 5. First, the full-rank training cost is calculated as
@@
where $E$ is the total number of epochs and $d$ is the model size. Then, the low-rank training cost is
@@
Thus, if $\rho$ is the same, $T_{DR}$ should be higher than $T_{SVD}$ and lower than $T_{F}$. E.g., ResNet20 contains 272,762 trainable parameters and the reparameterized model with SVD ($\rho=0.5$) has 155,170 parameters. When $E=150$, $I=60$, and $D=135$, where $\phi=0.5$, $T_{DR}/T_{F}$ becomes $0.7844$, which is shown in Comp. column in Table 4. Our empirical study shows that a relatively small DID-I, achieved by carefully tuning $I$ and $D$, yields accuracy comparable to full-rank training, while bringing the cost close to that of conventional low-rank training.
@@ 6 Conclusion
In this study, we propose a dynamic-rank training framework that restores the effective rank of model weights while preserving the efficiency of low-rank training. We also present two key insights for maximizing rank recovery by strategically interleaving full-rank epochs within low-rank training. Our extensive empirical study demonstrates that scheduling full-rank epochs at the end of the high-noise regime and the beginning of the low-noise regime maximizes the recovery of the model’s effective rank, thereby improving accuracy. The proposed dynamic-rank training scheme is general and readily applicable to any deep learning applications. We believe that harmonizing the proposed dynamic-rank training and other compute-efficient neural network training methods can be a promising future work. Additional discussion on potential limitations and future work is provided in Appendix 7.7.